Problem 26.79 Revieww Part A A small insect viewed through a convex lens is 2.3 cm from the lens and appears 2.5 times larger than its actual size What is the focal length of the lens? Express your answer using two significant figures. 四? cm SubmitPrev Previous Answers Request Answer
Answer #1

Solution)

Given,

Distance of the object, do = 2.3 cm

Magnification , m= 2.5

We know, Distance of image, di = - m do = - 5.75 cm

We know, 1 / f = ( 1/di ) + ( 1/ do )

1 / f = ( 1/- 5.75) + ( 1/ 2.3 )

Focal length, f =-3.83 cm (Ans)

=======